### v1

@@ 1 Introduction
Traditional efforts toward automating the extraction of structured data from clinical text have relied heavily on rule-based systems or shallow machine learning models (e.g., Conditional Random Fields, Support Vector Machines), each requiring extensive domain-specific feature engineering [uzuner2011, eftimov2017rule, ctakes, deepphe]. In the last decade, however, transformer-based models [vaswani2017], such as Bidirectional Encoder Representations from Transformers (BERT)[devlin2018bert] and its domain-specific variants like ClinicalBERT [alsentzer2019publicly, lee2020biobert, Preston2023Toward], have significantly improved upon classical approaches across a range of clinical natural language processing (NLP) tasks. More recently, the approach has shifted from fine-tuning domain specific models [lin2023oncobert, beltagy2019scibert] to using the generalized abilities of frontier large language models (LLMs) like GPT-4 and GPT-5[openai2023gpt4, openai2025gpt5] to extract key concepts from EHR records [Bhattarai2024GPT4].
Nevertheless, prior works often assume relatively uniform data (such as a single cancer diagnosis or one type of note), single-document inputs, or a very limited set of concepts. Published studies have demonstrated entity extraction for discrete variables such as tumor stage from pathology reports [abedian2021automated, kefeli2024tnm], receptor or genomic biomarkers from pathology or genomic reports [gauthier2022rwe, pironet2021receptor], and initial regimens or lines of therapy for select cancers [zeng2021initialtx, meng2021lot]. However, these approaches extract variables independently without synthesizing findings across multiple documents to resolve contradictions or complete partial information. This study is distinct in that it targets patient-level synthesis across heterogeneous, longitudinal records, where oncology-specific concepts must be inferred by collating and contextualizing information scattered across multiple notes and data types, rather than relying on single-step extraction from a uniform document.
@@ 2 Related Work
Early Clinical NLP and Rule-Based Methods: Early clinical NLP relied primarily on rule-based systems, utilizing domain-specific lexicons and hand-crafted patterns [friedman2004, chapman2001, ctakes]. These systems, though precise in certain scenarios, required substantial manual effort and were brittle when encountering variability [uzuner2012evaluating, uzuner2011]. Notable examples include MedXN for medication extraction [sohn2014medxn], comprehensive clinical NLP toolkits like CLAMP [soysal2018clamp], fracture identification from radiology reports [tibbo2019radiology], and sudden cardiac death risk factor extraction [moon2019extraction]. Statistical machine learning models, notably Conditional Random Fields (CRFs) and Support Vector Machines (SVMs), subsequently improved performance on clinical text tasks like de-identification and entity extraction [uzuner2011]. Early language modeling approaches showed promise for identifying relevant information in clinical notes [zhang2014automated], though still required task-specific adaptation. Nevertheless, adapting these models to heterogeneous oncology datasets posed significant challenges due to diverse terminologies and complex documentation [murff2011automated]. The evolution of clinical information extraction methods has been comprehensively reviewed by Wang et al. [wang2018clinical], documenting the systematic transition from rule-based to neural approaches.
@@
Deep Learning and Transformer Models in Clinical NLP: Transformer architectures, particularly BERT [vaswani2017, devlin2018bert], significantly advanced clinical NLP by providing powerful language representations. Domain-specific models such as BioBERT [lee2020biobert], ClinicalBERT [alsentzer2019publicly], and SciBERT [beltagy2019scibert] further enhanced clinical NLP tasks, achieving state-of-the-art performance in clinical entity recognition and relation extraction [peng2019transfer, si2019deep]. Comprehensive evaluations of these models on biomedical text-mining tasks demonstrated their superiority over traditional approaches [peng2019empirical]. However, oncology data, characterized by extensive and contextually complex documentation, poses unique challenges inadequately addressed by traditional single-pass transformers. Approaches employing hierarchical transformers and multi-document summarization have attempted to mitigate these limitations [ding2020cogran, fabbri2019multi, Preston2023Toward, Adamson2023Approach] but still fall short of fully replicating human abstraction.
@@
Large Language Models in Healthcare: Recent advances in large language models (LLMs), including GPT-4, and GPT-5[openai2023gpt4, openai2025gpt5], have demonstrated impressive capabilities in summarization, medical question answering, and clinical decision support [kung2023performance, agrawal2023large]. This has led to several research works exploring the capabilities of these models in extracting structured data from notes [Bhattarai2024GPT4, wong2025universal, porter2024llmd]. Recent multi-institutional efforts have shown promise in extracting social determinants of health from clinical notes using LLMs, achieving F1 scores over 0.9 [keloth2025social]. Work on extracting functional status information, including mobility assessments, from clinical notes has demonstrated the potential of LLMs for capturing complex clinical concepts [fu2024federated, kaster2025automated]. Despite this progress, single-pass LLM approaches struggle with multi-document EHR data, potentially losing critical context or misinterpreting conflicting information. Moreover, most research work so far has been limited to a limited set of oncology concepts [Adamson2023Approach, Preston2023Toward, wong2025universal, stuhlmiller2025scalable] instead of scaling it to a comprehensive RWD dataset dictionary.
@@
Agentic and Iterative NLP Frameworks: Recent NLP frameworks have increasingly adopted iterative reasoning approaches. The ReAct framework introduced by Yao et al. [yao2022react] enables LLMs to interleave reasoning and retrieval actions, improving accuracy on complex reasoning tasks. Techniques such as chain-of-thought prompting [wei2022chain] and self-consistency decoding [wang2022self] further enhance multi-step reasoning capabilities. Additionally, integrating external tools or retrieval systems, as explored in Toolformer [schick2023toolformer] and web-augmented retrieval methods [nakano2021webgpt], has shown promise. Yet, systematically integrating these approaches for clinical NLP, particularly oncology-specific data extraction, remains an open challenge due to accuracy and validation constraints.
@@
Autonomous & Agentic LLM Systems in Oncology: Oncology specific AI research has also begun moving from single-prompt LLM use toward agentic pipelines that interleave reasoning and tool calls. Ferber et al. developed an autonomous GPT-4–driven clinical-decision agent that orchestrates vision models, knowledge bases, and web search to solve multimodal oncology cases, raising accuracy from 30% (plain GPT-4) to over 87% on complex vignettes [ferber2025autonomous]. Sandhu et al. proposed an open-source modular framework that couples rule-based components with an agentic layer to generate comprehensive breast-cancer notes and benchmark treatment recommendations against NCCN guidelines [sandhu2025agentic]. Outside healthcare, Zhang and Elhamod proposed Data-to-Dashboard, a multi-agent architecture that detects domain context, extracts concepts, performs iterative self-reflection, and automatically builds analytic dashboards [zhang2025data_dashboard]. While these works demonstrate the promise of hierarchical agents, they tackle relatively narrow tasks (tens of test cases, disease-specific notes, or generic analytics) rather than end-to-end patient-level abstraction across thousands of heterogeneous documents.
@@
Domain-Specific LLMs for Oncology. Parallel efforts tailor foundation models and pipelines to oncology data. CancerBERT pretrains BERT on cancer corpora with a cancer-specific vocabulary and reports significant gains for extracting breast-cancer phenotypes (e.g., receptor status, site/laterality) over general and clinical BERT baselines from EHR notes and pathology reports[zhou2022cancerbert]. OncoBERT explores transfer learning on oncology notes for outcome prediction and phenotype structuring, reporting improvements on site and clinical T-staging tasks and highlighting interpretability considerations for radiation oncology workflows [lin2023oncobert, Preston2023Toward]. Beyond encoders, registry-oriented systems deliver patient-level abstraction: the original DeepPhe extracts phenotypes across entire EMRs, while DeepPhe-CR exposes API services integrated into cancer-registrar tools for computer-assisted abstraction, with usability studies and deployment guidance for registry workflows [deepphe, hochheiser2023deepphecr]. Patient-level staging at treatment initiation has also been demonstrated in Veterans Affairs data using rule-based NLP with validated roll-up for multiple myeloma [vaMMNLP2024]. In parallel, registry–EHR fusion cohorts constructed via NLP underscore the need for multi-source abstraction [ling2019mbc]. A complementary line targets narrow modalities or procedures—e.g., Woollie, a radiology-focused LLM trained on 39k impressions that attains strong progression-prediction AUROC of $0.97$ for progression prediction and outperforms general LLMs on medical benchmarks, and hybrid agentic pipelines coupling rules with GPT-4-Turbo for detailed spine-surgery variable extraction [zhu2025woollie, dagli2024ai_chart_review]. Orthogonal oncology IE efforts focus on single-entity extraction from pathology and clinical notes, including TNM staging and receptor status [abedian2021automated, kefeli2024tnm, pironet2021receptor], systemic treatment identification and line-of-therapy inference [zeng2021initialtx, meng2021lot], and broader RWD curation reviews and tooling [gauthier2022rwe, Zeng2019NLPPhenotyping].
@@
Most oncology LLM studies target narrow concept sets rather than comprehensive data dictionaries [Adamson2023Approach, Preston2023Toward, wong2025universal, stuhlmiller2025scalable]. These systems typically optimize for task- or modality-specific accuracy using single-pass encoders or rule-based components; they provide limited mechanisms for cross-entity dependency resolution, large-scale deduplication across documents, and curator-ready validation—gaps our agentic workflow aims to close.
@@
Table 1: Clinical Entities, Definitions, and Corresponding Attributes)
@@ 3.1 Problem Formulation
where $J$ is the total number of clinical entity types considered, and each $\hat{E}_{j}$ is a set of extracted instances of type $E_{j}$. Table 1 lists 21 oncology entity types considered in this study.
@@ 3.3.4 Pipeline Types
(a) Singe-Step Pipeline
@@ Acknowledgements
Hrituraj Singh and Yanshan Wang contributed to the conceptualization of the study. Methodology was developed by Hrituraj Singh, and Shashikant Gupta. Software and implementation were carried out by Shashikant Gupta, Jerrin John Thomas, and Arijeet Pramanik. Hrituraj Singh and Yanshan Wang were responsible for review and editing. All authors reviewed and approved the final manuscript.
@@
Hrituraj Singh, Shashikant Gupta, Arijeet Pramanik, and Jerrin John Thomas are employees of Triomics, Inc. Yanshan Wang declares no competing interests.

### v2

@@ 1 Introduction
Traditional efforts toward automating the extraction of structured data from clinical text have relied heavily on rule-based systems or shallow machine learning models (e.g., Conditional Random Fields, Support Vector Machines), each requiring extensive domain-specific feature engineering [1, 2, 3, 4]. In the last decade, however, transformer-based models [5], such as Bidirectional Encoder Representations from Transformers (BERT)[6] and its domain-specific variants like ClinicalBERT [7, 8, 9], have significantly improved upon classical approaches across a range of clinical natural language processing (NLP) tasks. More recently, the approach has shifted from fine-tuning domain specific models [10, 11] to using the generalized abilities of frontier large language models (LLMs) like GPT-4 and GPT-5[12, 13] to extract key concepts from EHR records [14].
Nevertheless, prior works often assume relatively uniform data (such as a single cancer diagnosis or one type of note), single-document inputs, or a very limited set of concepts. Published studies have demonstrated entity extraction for discrete variables such as tumor stage from pathology reports [15, 16], receptor or genomic biomarkers from pathology or genomic reports [17, 18], and initial regimens or lines of therapy for select cancers [19, 20]. However, these approaches extract variables independently without synthesizing findings across multiple documents to resolve contradictions or complete partial information. This study is distinct in that it targets patient-level synthesis across heterogeneous, longitudinal records, where oncology-specific concepts must be inferred by collating and contextualizing information scattered across multiple notes and data types, rather than relying on single-step extraction from a uniform document.
@@ 2 Related Work
Early Clinical NLP and Rule-Based Methods: Early clinical NLP relied primarily on rule-based systems, utilizing domain-specific lexicons and hand-crafted patterns [21, 22, 3]. These systems, though precise in certain scenarios, required substantial manual effort and were brittle when encountering variability [23, 1]. Notable examples include MedXN for medication extraction [24], comprehensive clinical NLP toolkits like CLAMP [25], fracture identification from radiology reports [26], and sudden cardiac death risk factor extraction [27]. Statistical machine learning models, notably Conditional Random Fields (CRFs) and Support Vector Machines (SVMs), subsequently improved performance on clinical text tasks like de-identification and entity extraction [1]. Early language modeling approaches showed promise for identifying relevant information in clinical notes [28], though still required task-specific adaptation. Nevertheless, adapting these models to heterogeneous oncology datasets posed significant challenges due to diverse terminologies and complex documentation [29]. The evolution of clinical information extraction methods has been comprehensively reviewed by Wang et al. [30], documenting the systematic transition from rule-based to neural approaches.
@@
Deep Learning and Transformer Models in Clinical NLP: Transformer architectures, particularly BERT [5, 6], significantly advanced clinical NLP by providing powerful language representations. Domain-specific models such as BioBERT [8], ClinicalBERT [7], and SciBERT [11] further enhanced clinical NLP tasks, achieving state-of-the-art performance in clinical entity recognition and relation extraction [31, 32]. Comprehensive evaluations of these models on biomedical text-mining tasks demonstrated their superiority over traditional approaches [33]. However, oncology data, characterized by extensive and contextually complex documentation, poses unique challenges inadequately addressed by traditional single-pass transformers. Approaches employing hierarchical transformers and multi-document summarization have attempted to mitigate these limitations [34, 35, 9, 36] but still fall short of fully replicating human abstraction.
@@
Large Language Models in Healthcare: Recent advances in large language models (LLMs), including GPT-4, and GPT-5[12, 13], have demonstrated impressive capabilities in summarization, medical question answering, and clinical decision support [37, 38]. This has led to several research works exploring the capabilities of these models in extracting structured data from notes [14, 39, 40]. Recent multi-institutional efforts have shown promise in extracting social determinants of health from clinical notes using LLMs, achieving F1 scores over 0.9 [41]. Work on extracting functional status information, including mobility assessments, from clinical notes has demonstrated the potential of LLMs for capturing complex clinical concepts [42, 43]. Despite this progress, single-pass LLM approaches struggle with multi-document EHR data, potentially losing critical context or misinterpreting conflicting information. Moreover, most research work so far has been limited to a limited set of oncology concepts [36, 9, 39, 44] instead of scaling it to a comprehensive RWD dataset dictionary.
@@
Agentic and Iterative NLP Frameworks: Recent NLP frameworks have increasingly adopted iterative reasoning approaches. The ReAct framework introduced by Yao et al. [45] enables LLMs to interleave reasoning and retrieval actions, improving accuracy on complex reasoning tasks. Techniques such as chain-of-thought prompting [46] and self-consistency decoding [47] further enhance multi-step reasoning capabilities. Additionally, integrating external tools or retrieval systems, as explored in Toolformer [48] and web-augmented retrieval methods [49], has shown promise. Yet, systematically integrating these approaches for clinical NLP, particularly oncology-specific data extraction, remains an open challenge due to accuracy and validation constraints.
@@
Autonomous & Agentic LLM Systems in Oncology: Oncology specific AI research has also begun moving from single-prompt LLM use toward agentic pipelines that interleave reasoning and tool calls. Ferber et al. developed an autonomous GPT-4–driven clinical-decision agent that orchestrates vision models, knowledge bases, and web search to solve multimodal oncology cases, raising accuracy from 30% (plain GPT-4) to over 87% on complex vignettes [50]. Sandhu et al. proposed an open-source modular framework that couples rule-based components with an agentic layer to generate comprehensive breast-cancer notes and benchmark treatment recommendations against NCCN guidelines [51]. Outside healthcare, Zhang and Elhamod proposed Data-to-Dashboard, a multi-agent architecture that detects domain context, extracts concepts, performs iterative self-reflection, and automatically builds analytic dashboards [52]. While these works demonstrate the promise of hierarchical agents, they tackle relatively narrow tasks (tens of test cases, disease-specific notes, or generic analytics) rather than end-to-end patient-level abstraction across thousands of heterogeneous documents.
@@
Domain-Specific LLMs for Oncology. Parallel efforts tailor foundation models and pipelines to oncology data. CancerBERT pretrains BERT on cancer corpora with a cancer-specific vocabulary and reports significant gains for extracting breast-cancer phenotypes (e.g., receptor status, site/laterality) over general and clinical BERT baselines from EHR notes and pathology reports[53]. OncoBERT explores transfer learning on oncology notes for outcome prediction and phenotype structuring, reporting improvements on site and clinical T-staging tasks and highlighting interpretability considerations for radiation oncology workflows [10, 9]. Beyond encoders, registry-oriented systems deliver patient-level abstraction: the original DeepPhe extracts phenotypes across entire EMRs, while DeepPhe-CR exposes API services integrated into cancer-registrar tools for computer-assisted abstraction, with usability studies and deployment guidance for registry workflows [4, 54]. Patient-level staging at treatment initiation has also been demonstrated in Veterans Affairs data using rule-based NLP with validated roll-up for multiple myeloma [55]. In parallel, registry–EHR fusion cohorts constructed via NLP underscore the need for multi-source abstraction [56]. A complementary line targets narrow modalities or procedures—e.g., Woollie, a radiology-focused LLM trained on 39k impressions that attains strong progression-prediction AUROC of $0.97$ for progression prediction and outperforms general LLMs on medical benchmarks, and hybrid agentic pipelines coupling rules with GPT-4-Turbo for detailed spine-surgery variable extraction [57, 58]. Orthogonal oncology IE efforts focus on single-entity extraction from pathology and clinical notes, including TNM staging and receptor status [15, 16, 18], systemic treatment identification and line-of-therapy inference [19, 20], and broader RWD curation reviews and tooling [17, 59].
@@
Most oncology LLM studies target narrow concept sets rather than comprehensive data dictionaries [36, 9, 39, 44]. These systems typically optimize for task- or modality-specific accuracy using single-pass encoders or rule-based components; they provide limited mechanisms for cross-entity dependency resolution, large-scale deduplication across documents, and curator-ready validation—gaps our agentic workflow aims to close.
@@
Table 1: Clinical Entities, Definitions, and Corresponding Attributes
@@ 3.1 Problem Formulation
where $J$ is the total number of clinical entity types considered, and each $\hat{E}_{j}$ is a set of extracted instances of type $E_{j}$. Table 1 lists 16 oncology entity types considered in this study.
@@ 3.3.4 Pipeline Types
(a) Single-Step Pipeline
@@ Acknowledgements
Hrituraj Singh and Yanshan Wang contributed to the conceptualization of the study. Methodology was developed by Hrituraj Singh, and Shashi kant Gupta. Software and implementation were carried out by Shashikant Gupta, Jerrin John Thomas, and Arijeet Pramanik.. All authors reviewed and approved the final manuscript.
@@
Hrituraj Singh, Shashikant Gupta, Arijeet Pramanik, Regina Schwind and Jerrin John Thomas are employees of Triomics, Inc.
Lauren Wiener, Avi Raju, Jeremy Kornbluth, and Zhaohui Su are employees of Ontada LLC.
Yanshan Wang declares no competing interests.